# Efficient Post-Training Refinement of Latent Reasoning in Large Language Models

### v1

@@ 5.1 Chain of Thought Reasoning
Chain-of-Thought (CoT) prompting enhances reasoning in LLMs by decomposing complex problems into step-by-step textual reasoning traces Wei et al. (2022b); Kojima et al. (2022). Later works improve CoT with better aggregation (e.g., self-consistency Wang et al. (2022b)), scalable prompt generation Zhang et al. (2022), and tree-structured exploration like Tree-of-Thoughts Yao et al. (2023a). Prompting strategies have also been optimized through complexity-aware design Fu et al. (2022) and iterative demonstration Nye et al. (2021). These techniques have been widely applied to domains such as math Zhou et al. (2022), commonsense Huang et al. (2022), and symbolic logic Liu et al. (2023), showing broad improvements. Extensions such as Selection-Inference Creswell et al. (2022), Program-of-Thoughts Chen et al. (2022), and Multimodal-CoT Lu et al. (2022) further combine CoT with selection mechanisms, symbolic programs, or visual reasoning inputs.
5.2 Latent-Space Reasoning in LLMs
@@
Recent work proposes moving reasoning from token-level generation to latent space updates, enabling more compact and abstract reasoning. Coconut Hao et al. (2024) is a foundational method that replaces intermediate token outputs with internal latent states that evolve step by step. Other approaches expand this idea through latent sampling (LaTRO Chen et al. (2024)), self-training to uncover latent reasoning (SERT Zhang et al. (2025)), or expectation-maximization loops for iterative reasoning Ruan et al. (2025). Looped transformers Saunshi et al. (2025) reuse a smaller model multiple times to simulate deep reasoning trajectories. In applied domains, ReaRec Tang et al. (2025) introduces latent multi-step reasoning into recommender systems, demonstrating the broader utility of latent-state methods beyond language tasks.
6 Conclusion Remarks
@@
We present a latent reasoning framework for LLM by introducing residual refinement and contrastive latent search to improve stability and enable dynamic self-correction without retraining. The new framework operates in a training-free, plug-and-play manner. Experiments show that our method improves accuracy by 2–5% over latent-only reasoning across five benchmarks, with up to 7.7% gain on ProsQA. The improvement demonstrates the effectiveness of latent refinement with internal information. We show that even without token-level outputs or supervised feedback, LLMs can adjust internal belief states using post-hoc representational updates, pointing toward a new direction of self-corrective, embedding-level reasoning.
Limitations
While our framework improves reasoning performance across multiple tasks and models, it still has several limitations. (1) Even though the method can be applied to almost all open-source LLMs, we just test it on some relatively small LLMs. So the results are not comparable to the latest model, like ChatGPT-4.1 or DeepSeek-R1. In addition, we do not compare with sampling-based or search-based CoT variants (e.g., Tree-of-Thought Yao et al. (2023a), Self-Consistency Wang et al. (2022b)), since our focus is on deterministic, internal latent refinement without additional decoding or re-ranking. (2) Latent reasoning may not be optimal for all types of tasks. For example, human tend to solve math problems with written-down logic, which involves explicit reasoning processes. In the future study, the routing method Wang et al. (2025b) can be applied to dynamically reason the query. (3) Latent reasoning lacks human-readable intermediate steps, making it harder to interpret or debug compared to token-based methods. (4) Inspired by LLM agent researches, the future work can introduces tools to ensure stronger, more correct, and more reasonable reasoning processes.
@@ Appendix A Algorithm Overview
Here is the Algorithm pseudo-code of our train-free framework.
@@
Appendix B Dataset Descriptions
GSM8K Cobbe et al. (2021) [^1] https://huggingface.co/datasets/openai/gsm8k: A dataset of 8.5K high-quality grade school math word problems created by human problem writers. It is divided into 7.5K training and 1K test examples. Each problem requires 2–8 steps of basic arithmetic reasoning (addition, subtraction, multiplication, division) and includes a detailed natural language solution explanation.
MathQA Amini et al. (2019) [^2] https://huggingface.co/datasets/allenai/math_qa: A dataset containing approximately 37.2K math word problems derived from AQuA. Each problem is annotated with a programmatic solution template using a domain-specific language comprising 58 operators. The dataset covers various mathematical domains, including percentages, geometry, and linear equations, and provides multiple-choice answers.
AQUA-RAT Ling et al. (2017) [^3] https://huggingface.co/datasets/deepmind/aqua_rat: A large-scale dataset consisting of approximately 100K algebraic word problems. Each question is accompanied by a step-by-step natural language rationale explaining the solution process. The dataset is designed to train models capable of generating both the solution and the explanatory rationale.
StrategyQA Geva et al. (2021) [^4] https://huggingface.co/datasets/voidful/StrategyQA: A question-answering benchmark comprising 2,780 examples. Each question requires implicit multi-hop reasoning, where the necessary reasoning steps are not explicitly stated and must be inferred. The dataset includes decompositions and evidence paragraphs for each question.
ProsQA Hao et al. (2024): A diagnostic dataset designed to evaluate models’ planning abilities. It features examples that require multi-step reasoning and the ability to handle distractors. Each question is paired with a human-written rationale, facilitating the assessment of models’ reasoning processes.
Appendix C Chain-of-Thought Data Construction
@@
  • Latent CoT (Latent Reasoning): We follow the Coconut Hao et al. (2024) setup. The model is trained to generate the answer $y$ from input $x$, but the reasoning steps $s$ are used as supervision in the latent space. They are not generated as tokens, but guide the intermediate hidden representations.
@@
Appendix D Implementation and Hyperparameters
@@
Appendix E Experimental Environment
@@
Appendix F Generalization to Unseen Tasks
@@
Model Trained on MathQA Trained on GSM8K Difference
@@
These results highlight the challenge of cross-task generalization in small and medium-sized LLMs. Despite using latent representations and post-training refinement, our method—like CoT and Coconut—fails to transfer reasoning skills between datasets. We believe this is due to two factors: (1) models like GPT-2 lack the scale and capacity needed for emergent generalization Wei et al. (2022a); (2) our method is designed to refine reasoning trajectories within a task, not to perform domain adaptation. While effective in-distribution, the framework does not substitute for task-specific training when reasoning patterns differ across domains.
Appendix G More Output Case Studies
@@ G.3 Example 3
Appendix H Broader Impact
The proposed post-training framework to refine the latent reasoning trajectory can also be applied to many downstream applications such as data-centric AI Wang et al. (2025a); Ying et al. (2025b, 2024b, 2024c, 2024d, 2023, 2024a); Bai et al. ; Gong et al. (2025a, c, b, e, d); Wang et al. (2024b), linguistics Ying et al. (2020); Wang et al. (2022a), business (He et al., 2024; Wang et al., 2024c; Bai et al., 2023, 2024; Li et al., 2023), computer vision Hu et al. (2022); Dong et al. (2023), and medicine Liu et al. (2019); Wang et al. (2022c); Liu et al. (2024b); Wang et al. (2024a); Liu et al. (2024a); Li et al. (2024). We can also apply the routing strategy Wang et al. (2025b) to handle situations with multiple reasoning strategies or LLMs. The reinforcement learning strategy Ying et al. (2025a) and rule-based augmentation Bai et al. (2025) could also be applied to enhance the robustness of the reasoning process.

### v2

@@ Introduction
Chain-of-Thought (CoT) prompting enhances reasoning in LLMs by decomposing complex problems into step-by-step textual reasoning traces (wei2022chain; kojima2022large). Later works improve CoT with better aggregation (e.g., self-consistency (wang2022self)), scalable prompt generation (zhang2022automatic), and tree-structured exploration like Tree-of-Thoughts (yao2023tree). Prompting strategies have also been optimized through complexity-aware design (fu2022complexity) and iterative demonstration (nye2021show). These techniques have been widely applied to domains such as math (zhou2022least), commonsense (huang2022language), and symbolic logic (liu2023glore), showing broad improvements. Extensions such as Selection-Inference (creswell2022selection), Program-of-Thoughts (chen2022program), and Multimodal-CoT (lu2022learn) further combine CoT with selection mechanisms, symbolic programs, or visual reasoning inputs.
Latent-Space Reasoning in LLMs
@@
Recent work proposes moving reasoning from token-level generation to latent space updates, enabling more compact and abstract reasoning. Coconut (hao2024training) is a foundational method that replaces intermediate token outputs with internal latent states that evolve step by step. Other approaches expand this idea through latent sampling (LaTRO (chen2024language)), self-training to uncover latent reasoning (SERT (zhang2025self)), or expectation-maximization loops for iterative reasoning (ruan2025reasoning). Looped transformers (saunshi2025reasoning) reuse a smaller model multiple times to simulate deep reasoning trajectories. In applied domains, ReaRec (tang2025think) introduces latent multi-step reasoning into recommender systems, demonstrating the broader utility of latent-state methods beyond language tasks.
Conclusion Remarks
@@
We present a latent reasoning framework for LLM by introducing residual refinement and contrastive latent search to improve stability and enable dynamic self-correction without retraining. The new framework operates in a training-free, plug-and-play manner. Experiments show that our method improves accuracy by 2–5% over latent-only reasoning across five benchmarks, with up to 7.7% gain on ProsQA. The improvement demonstrates the effectiveness of latent refinement with internal information.
Our method is simple, efficient, and general. It works across different model architectures and sizes, and adds minimal overhead during inference. Even without generating intermediate text or using extra supervision, LLMs can still adjust their internal states through small updates in the latent space. This shows a new way to make models correct themselves at the embedding level, which can be useful for frozen models or low-resource settings.
Acknowledgments
Dr. Yanjie Fu is supported by the National Science Foundation (NSF) via the grant numbers: 2426340, 2416727, 2421865, 2421803. Dr. Kunpeng Liu is supported by the National Science Foundation (NSF) via the grant numbers 2550105, 2550106, and 2242812.
@@ Appendix A Algorithm Overview
This is the pseudo-code of our train-free framework.
@@
Appendix B Code Repository
Our method is based on Coconut (hao2024training). Coconut codes are available at: https://github.com/facebookresearch/coconut.
The codes and data used in this paper are publicly available at this anonymous repository: https://github.com/anord-wang/Lateng-Reasoning.
Appendix C Dataset Descriptions
GSM8K (cobbe2021training) [^1] https://huggingface.co/datasets/openai/gsm8k: A dataset of 8.5K high-quality grade school math word problems created by human problem writers. It is divided into 7.5K training and 1K test examples. Each problem requires 2–8 steps of basic arithmetic reasoning (addition, subtraction, multiplication, division) and includes a detailed natural language solution explanation.
MathQA (amini2019mathqa) [^2] https://huggingface.co/datasets/allenai/math˙qa: A dataset containing approximately 37.2K math word problems derived from AQuA. Each problem is annotated with a programmatic solution template using a domain-specific language comprising 58 operators. The dataset covers various mathematical domains, including percentages, geometry, and linear equations, and provides multiple-choice answers.
AQUA-RAT (ling2017program) [^3] https://huggingface.co/datasets/deepmind/aqua˙rat: A large-scale dataset consisting of approximately 100K algebraic word problems. Each question is accompanied by a step-by-step natural language rationale explaining the solution process. The dataset is designed to train models capable of generating both the solution and the explanatory rationale.
StrategyQA (geva2021did) [^4] https://huggingface.co/datasets/voidful/StrategyQA: A question-answering benchmark comprising 2,780 examples. Each question requires implicit multi-hop reasoning, where the necessary reasoning steps are not explicitly stated and must be inferred. The dataset includes decompositions and evidence paragraphs for each question.
ProsQA (hao2024training): A diagnostic dataset designed to evaluate models’ planning abilities. It features examples that require multi-step reasoning and the ability to handle distractors. Each question is paired with a human-written rationale, facilitating the assessment of models’ reasoning processes.
Appendix D Chain-of-Thought Data Construction
@@
  • Latent CoT (Latent Reasoning): We follow the Coconut (hao2024training) setup. The model is trained to generate the answer $y$ from input $x$, but the reasoning steps $s$ are used as supervision in the latent space. They are not generated as tokens, but guide the intermediate hidden representations.
@@
Appendix E Implementation and Hyperparameters
@@
Appendix F Experimental Environment
@@
Appendix G Generalization to Unseen Tasks
@@
Model
Trained on
MathQA
Trained on
GSM8K
Difference
@@
These results highlight the challenge of cross-task generalization in small and medium-sized LLMs. Despite using latent representations and post-training refinement, our method—like CoT and Coconut—fails to transfer reasoning skills between datasets. We believe this is due to two factors: (1) models like GPT-2 lack the scale and capacity needed for emergent generalization (wei2022emergent); (2) our method is designed to refine reasoning trajectories within a task, not to perform domain adaptation. While effective in-distribution, the framework does not substitute for task-specific training when reasoning patterns differ across domains.
Appendix H More Output Case Studies
@@
Appendix I Limitations
While our framework improves reasoning performance across multiple tasks and models, it still has several limitations. (1) Even though the method can be applied to almost all open-source LLMs, we just test it on some relatively small LLMs. So the results are not comparable to the latest model, like ChatGPT-4.1 or DeepSeek-R1. In addition, we do not compare with sampling-based or search-based CoT variants (e.g., Tree-of-Thought (yao2023tree), Self-Consistency (wang2022self)), since our focus is on deterministic, internal latent refinement without additional decoding or re-ranking. (2) Latent reasoning may not be optimal for all types of tasks. For example, human tend to solve math problems with written-down logic, which involves explicit reasoning processes. In the future study, the routing method (wang2025mixllm) can be applied to dynamically reason the query. (3) Latent reasoning lacks human-readable intermediate steps, making it harder to interpret or debug compared to token-based methods. (4) Inspired by the research of LLM agents (wang2025dataforge), future work can introduce tools to ensure stronger, more correct, and more reasonable reasoning processes.
Appendix J Broader Impact
The proposed post-training framework to refine the latent reasoning trajectory can also be applied to many downstream applications such as data-centric AI (wang2025towards; ying2025survey; ying2024feature; ying2024revolutionizing; ying2024unsupervised; ying2023self; ying2024topology; baiprivacy; gong2025agentic; gong2025evolutionary; gong2025sculpting; gong2025neuro; gong2025unsupervised; wang2024knockoff), linguistics (ying2020sichuan; wang2022hierarchal), business (rs1; wang2024llm; rs2; rs3; li2023sehf), computer vision (hu2022transrac; dong2023weakly; dong2025mmtok), and medicine (liu2019edta; wang2022successful; liu2024calorie; wang2024lcmdc; liu2024pth; li2024sade). We can also apply the routing strategy (wang2025mixllm) to handle situations with multiple reasoning strategies or LLMs. The reinforcement learning strategy (ying2025bridging) and rule-based augmentation (bai2025brownian) could also be applied to enhance the robustness of the reasoning process.